Consider the following Gibbs energies at 25 °C. Substance AGOf (kJ mo Ag (aq) 77.1 Cl (aq) 131.2 109.8 AgCl(s) 104.0 Br (aq) 96.9 AgBr(s) (b) Calculate the solubility-product constant of AgCl. (a) Calculate AGO for the dissolution of AgCl(s) rXn Number Number K- O kJ mol (c) Calculate AGO for the dissolution of AgBr(s) (d) Calculate the solubility-product constant of AgBr rXn Number Number K- kJ molPlease show work
Answer #1

a) AgCl \rightarrow Ag+  + Cl-

\DeltaGrxn0  = 77.1-131.2 + 109.8 = 55.7

b) \DeltaG = \DeltaG0  + RTln(K)

\DeltaG = 0 at the equlribium

K = exp(-\DeltaG0/RT)

K = 1.726 x 10-10

c) AgBr \rightarrow Ag+ + Br-

\DeltaGrxn0= 77.1-104 +96.9 = 70

d) \DeltaG = \DeltaG0  + RTln(K)

\DeltaG = 0 at the equlribium

K = exp(-\DeltaG0/RT)

K = 5.384 x 10-13